# Text2SQL-Flow: A Robust SQL-Aware Data Augmentation Framework for Text-to-SQL

### v2

@@ IV-A1 Benchmarks
  • Spider [yu2018spider] contains 10,181 NL questions and 5,693 unique SQL queries, involving 200 databases across multiple tables covering 138 different domains. We utilize both its development set, containing 1,034 samples, and its test set with 2,147 samples for evaluation.
  • BIRD [li2024can] includes 12,751 NL question-SQL pairs and 95 databases across 37 domains, focusing on large-scale noisy data and external knowledge reasoning. Our experiments are performed on its development set, which consists of 1,534 data samples.
  • EHRSQL [lee2022ehrsql] includes approximately 24,000 NL question-SQL pairs linked to 2 open-source electronic health record databases. The dataset introduces challenges such as complex, time-sensitive questions and the inclusion of unanswerable questions.
  • Spider-DK, Spider-Syn, and Spider-Realistic [gan2021exploring, gan2021towards, deng2020structure] are three widely-adopted robustness benchmarks designed to evaluate model performance in practical Text-to-SQL applications. Spider-DK tests the model’s ability to understand implicit domain knowledge in NL questions, offering 535 samples for evaluation. Spider-Syn replaces schema-related words with manually selected synonyms and provides 1,034 evaluation samples. Spider-Realistic removes or paraphrases explicit mentions of column names, offering 508 samples for evaluation.
IV-A2 Baselines
Primarily, we evaluate the model trained on our SQLFlow dataset against a range of both open-source and closed-source mainstream LLMs. The closed-source models include GPT-4o-mini [openai2024gpt4omini], GPT-4o [hurst2024gpt], and GPT-4-Turbo [achiam2023gpt]. The open-source models encompass DeepSeek-Coder-7B-Instruct [guo2024deepseek], Qwen2.5-Coder-7B-Instruct [hui2024qwen2], Qwen2.5-7B-Instruct [qwen2025qwen25], OpenCoder-8B-Instruct [huang2024opencoder], Meta-Llama-3.1-8B-Instruct [grattafiori2024llama], Granite-8B-Code-Instruct [mishra2024granite], and Granite-3.1-8B-Instruct [ibm2024granite31]. All these open-source models have parameter counts of either 7B or 8B, matching the scale of our trained model. As for the training data of the models, we utilize not only the public datasets Spider-train [yu2018spider] and BIRD-train [li2024can], but also the synthetic dataset SynSQL proposed by OmniSQL [li2025omnisql]. To increase the diversity of our comparisons, for both SynSQL and SQLFlow, regardless of their original data sources, we randomly sample two subsets: one containing 50K examples, which constitutes a suitably large scale for comparison, and another containing 90K examples (precisely 89,544 examples), matching the exact size of the SQLFlow dataset for ease of reference. Additionally, we also use the full SynSQL-2.5M dataset, comprising 2.5 million examples, as a reference.
Secondly, for the few-shot example retrieval, we compare our retrieval method with classic example retrieval strategies. All similarity calculations use cosine similarity.
  • Zero-shot [kojima2022large] refers to a setting in which no few-shot examples are used during inference.
  • Random Selection means selecting the few-shot examples from the knowledge base randomly without any criteria.
  • Question Similarity Selection [liu2023comprehensive] identifies relevant examples by measuring the similarity between NL questions. It calculates the similarity between the target question and all candidate questions and selects those with the highest similarity.
[…]
  • DAIL Selection [gao2023text] retrieves examples by considering similarity in both the question and query spaces. It begins by masking domain-specific words in the target and candidate questions. The candidates are then ranked by question similarity. This ranking is subsequently filtered, retaining only examples whose query Jaccard similarity to the target query exceeds a predefined threshold. Since the question initially lacks a corresponding SQL, a SQL is pre-generated to serve as a reference.
TABLE II: Performance of LLMs on mainstream benchmarks. The first two blocks list closed-source and open-source base models, respectively. The last two blocks show fine-tuned models, where the first column indicates the training data setting.
@@ IV-A2 Baselines
SynSQL(50k) 67.1 73.9 72.7 78.6 49.1 55.2 33.6 40.8 63.8 66.1 59.6 63.5 69.3 71.6 59.3 64.2
SynSQL(90k) 68.2 74.6 73.4 78.5 51.1 54.9 31.8 38.0 61.8 67.4 58.9 63.6 69.0 70.9 59.2 64.0
SynSQL(2.5M) 70.6 73.7 78.3 82.5 58.9 62.0 35.1 37.0 72.3 74.7 61.0 63.1 67.9 69.4 63.4 66.1
Spider+BIRD+SQLFlow 74.9 79.2 78.4 82.3 53.4 58.9 28.4 36.5 67.7 69.7 66.6 69.1 74.4 75.0 63.4 67.2
SQLFlow(50k) 69.9 76.8 75.1 80.1 51.4 57.6 28.0 36.4 65.9 68.1 61.3 67.5 69.6 73.5 60.2 65.7
SQLFlow 71.4 76.4 75.8 80.0 54.6 56.8 55.5 56.3 66.5 67.7 61.6 67.3 71.4 72.7 65.3 68.2
@@
Spider+BIRD+SQLFlow 85.5 87.5 87.5 88.5 58.3 64.0 27.9 39.8 71.0 73.1 75.0 76.2 82.3 83.7 69.6 73.3
SQLFlow(50k) 80.9 84.9 84.6 85.8 57.9 62.5 27.8 39.4 69.7 71.2 70.0 74.0 77.8 82.1 67.0 71.4
SQLFlow 82.0 85.0 84.8 86.0 59.2 61.5 56.1 58.7 69.7 71.0 69.9 74.4 79.5 81.7 71.6 74.0
IV-A3 Evaluation Metric
@@
IV-A4 Implementation Details
@@
To validate the effectiveness of our framework, we conduct full-parameter SFT on two widely adopted open-source LLMs: Meta-Llama-3.1-8B-Instruct [grattafiori2024llama] and Qwen2.5-Coder-7B-Instruct [hui2024qwen2]. The training configuration includes a learning rate of $2\times 10^{-5}$, 2 training epochs, and a warmup ratio of 15%. During LLM inference, we explore two decoding strategies: greedy decoding and majority voting. Greedy decoding (denoted as Gre in tables) uses a temperature of 0 to produce deterministic outputs. Majority voting (denoted as Maj in tables) samples 8 candidate responses per input at a temperature of 0.8. From these candidates, valid SQL queries are extracted and executed; the final prediction is selected as the query whose execution result receives the most votes.
@@
As for the retrieval module, we develop an embedding model by full-parameter SFT a Qwen3-Embedding-0.6B model [zhang2025qwen3]. The model is trained for 3 epochs with a learning rate of $6\times 10^{-6}$, and a per-device batch size of 2. During training, we treat the NL question as the query and its corresponding SQL as the positive example. For negative examples, we randomly sample 30 SQL queries from the rest of the dataset. For inference, we compute Question-SQL similarity by generating embeddings with both the original and our fine-tuned embedding model, using cosine similarity. We apply the masking strategy of DAIL [gao2023text] during inference and training. A 5-shot setting is used during LLM inference. We use the GPT-4o as the closed-source LLM during the experiment with greedy decoding and set the temperature to zero. We follow the DAIL organization strategy when organizing the prompts.
@@
TABLE III: Performance comparison of retrieval strategies. “Upper limit”: retrieval using ground-truth SQL. “Mask”: whether query/example masking is applied. “API”: whether LLM API cost is incurred during retrieval.
Retrieval Strategy Mask API BIRD (dev) Spider (dev) Spider (test)
Baseline reference strategies
Zero-shot - - 52.5 74.1 75.5
Random selection 56.6 78.3 79.8
Question similarity 58.3 79.6 81.9
59.0 81.0 83.1
DAIL selection 58.8 80.0 82.2
59.4 80.8 82.9
Question-SQL similarity 58.7 80.9 81.7
Upper limit 61.8 83.9 85.7
Question-SQL similarity trained on Qwen3-embedding-0.6B
SQLFlow-part+Spider+BIRD 59.3 81.0 82.1
61.0 81.2 83.5
IV-A5 Training Data
For open-source LLMs, we use the entire SQLFlow dataset as training data. To train the embedding model, we use a filtered subset of the SQLFlow-Spider and SQLFlow-BIRD datasets (described in Section III-A8), retaining only SQL queries that begin with a SELECT statement. This yields 67,570 samples, which we refer to as SQLFlow-part. We also establish two baseline datasets for comparison. The first combines the standard Spider-train and BIRD-train benchmarks. The second, denoted SynSQL-part, is a random sample of 67,570 queries from SynSQL-2.5M [li2025omnisql], matched in size to SQLFlow-part for fair comparison. Table IV summarizes the used training datasets.
TABLE IV: Explanation of training data used.
Data Name Description Data Size
Spider+BIRD Training sets from Spider and BIRD. 18,087
SQLFlow Combination of SQLFlow-Spider, SQLFlow-BIRD, and SQLFlow-EHRSQL 89,544
SQLFlow-part Filtered subset retaining only SELECT-initiated SQL queries from SQLFlow-Spider and SQLFlow-BIRD 67,570
SynSQL-part Randomly sampled subset from SynSQL-2.5M, maintaining the same data volume as SQLFlow-part. 67,570
@@ IV-B Performance of LLM Fine-tuning
As shown in Table II, the data generated by our framework leads to consistent performance improvements across multiple mainstream benchmarks, demonstrating the effectiveness of Text2SQL-Flow. We fine-tune two widely used open-source models, Meta-Llama-3.1-8B-Instruct and Qwen2.5-Coder-7B-Instruct, to evaluate the general applicability of our data. In both cases, models trained on our generated data significantly outperformed their respective baselines as well as other models. However, due to inherent differences in model capabilities, Meta-Llama-3.1-8B-Instruct performed slightly worse than Qwen2.5-Coder-7B-Instruct. When fine-tuned with SQLFlow data, Qwen2.5-Coder-7B-Instruct achieved notable gains: execution accuracy (“Gre”) on Spider-dev increased from 73.4% to 82.0% (+8.6%), on BIRD-dev from 50.9% to 59.2% (+8.3%), and on the challenging EHRSQL benchmark from 24.3% to 56.1% (+21.8%). These results confirm two key points: (1) SFT substantially enhances model performance on Text-to-SQL tasks, and (2) the data generated via our Text2SQL-Flow framework exhibits high quality and strong training utility. The consistent improvements across two distinct model architectures demonstrate the broad effectiveness of our dataset in boosting model performance.
@@
In comparison with other training datasets, our data also demonstrates clear advantages. On comparable data scales (50K and 90K), our approach further exhibits superior performance compared to SynSQL. Specifically, on the Spider-test and BIRD-dev datasets, the model trained with SQLFlow(50K) achieves 84.6% and 57.9% execution accuracy (“Gre”), outperforming SynSQL(50K), which reaches 81.8% and 54.0%, respectively. These improvements highlight the higher quality and stronger generalization capabilities of our augmented training data. Likewise, the model trained with SQLFlow(90K) not only surpasses the baseline models but also outperforms SynSQL(90K), reinforcing the robustness of our data generation approach. Remarkably, even when trained on a much smaller dataset, the model fine-tuned with SQLFlow(90K) achieves performance comparable to SynSQL-2.5M on several challenging benchmarks. When incorporating real-world datasets such as Spider-train and BIRD-tain into the training data, the model achieves SOTA results across multiple benchmarks. We attribute these gains to the close alignment between the benchmark test distributions and our generated data, which enables more effective adaptation and leads to substantial performance improvements.
IV-C Comparison of Retrieval Strategies
We conduct a performance comparison between our proposed retrieval method and several baseline strategies, with results shown in Table III. The embedding model is trained on SQLFlow-part, Spider-train, and BIRD-train, as described in Table IV. The base model used in the experiments is Qwen3-Embedding-0.6B.

### v3

@@ IV-A1 Benchmarks
  • Spider [52] contains 10,181 NL questions and 5,693 unique SQL queries, involving 200 databases across multiple tables covering 138 different domains. We utilize both its development set, containing 1,034 samples, and its test set with 2,147 samples for evaluation.
  • BIRD [32] includes 12,751 NL question-SQL pairs and 95 databases across 37 domains, focusing on large-scale noisy data and external knowledge reasoning. Our experiments are performed on its development set, which consists of 1,534 data samples.
  • EHRSQL [26] includes approximately 24,000 NL question-SQL pairs linked to 2 open-source electronic health record databases. The dataset introduces challenges such as complex, time-sensitive questions and the inclusion of unanswerable questions.
  • Spider-DK, Spider-Syn, and Spider-Realistic [9, 8, 4] are three widely-adopted robustness benchmarks designed to evaluate model performance in practical Text-to-SQL applications. Spider-DK tests the model’s ability to understand implicit domain knowledge in NL questions, offering 535 samples for evaluation. Spider-Syn replaces schema-related words with manually selected synonyms and provides 1,034 evaluation samples. Spider-Realistic removes or paraphrases explicit mentions of column names, offering 508 samples for evaluation.
TABLE II: Performance of LLMs on mainstream benchmarks. The first two blocks list closed-source and open-source base models, respectively, which are included as reference baselines. The last two blocks present fine-tuned models, where the first column indicates the training data setting.“Seed Data” denotes the union of Spider-train, BIRD-train, and EHRSQL-train.
@@
SynSQL(75k) 78.1 80.9 78.1 81.8 47.7 54.9 28.9 35.9 66.4 67.5 67.5 72.1 74.0 77.0 63.0 67.2
Seed Data + SQLFlow 81.4 85.4 81.9 85.2 54.5 59.5 61.2 61.5 67.5 72.7 71.2 76.5 78.0 80.1 70.8 74.4
SQLFlow 80.4 83.3 81.3 83.9 51.0 57.1 51.2 56.0 66.0 72.0 69.3 74.0 75.4 79.7 67.8 72.3
@@
Seed Data + SQLFlow 84.9 86.2 85.3 87.3 58.0 62.1 61.2 62.0 74.4 75.9 74.8 78.0 82.1 83.5 74.4 76.4
SQLFlow 83.4 85.0 85.1 86.6 56.8 61.0 53.7 57.5 73.3 74.8 73.4 76.4 81.9 81.7 72.5 74.7
IV-A2 Evaluation Metric
@@
IV-A3 Baselines of LLM Training
Multiple open-source and closed-source mainstream LLMs are evaluated as reference baselines. The closed-source models include GPT-4-Turbo [1] and GPT-4o [21]. The open-source models encompass DeepSeek-Coder-7B-Instruct [16], Qwen2.5-Coder-7B-Instruct [20], OpenCoder-8B-Instruct [19], Meta-Llama-3.1-8B-Instruct [13], and Granite-8B-Code-Instruct [35]. Regarding the training data for the models, we utilize not only the public datasets Spider-train [52] and BIRD-train [32], but also the synthetic dataset SynSQL proposed by OmniSQL [29]. We randomly sample a subset of SynSQL containing 75K examples (75,386 instances), matching the exact size of the SQLFlow dataset for ease of comparison. In addition, we use the full SynSQL-2.5M dataset, comprising 2.5 million examples, as a reference. We use GPT-4o as the backend model to construct SQLFlow and set the temperature to zero.
IV-A4 Baselines of Few-shot Example Retrieval
We compare our retrieval method with multiple example retrieval strategies under both masked and unmasked settings. The masking strategy in DAIL-SQL [10] is applied here. All similarity calculations use cosine similarity. Strategies are as follows:
  • Zero-shot [24] refers to a setting in which no few-shot examples are used during inference.
  • Random Selection means selecting the examples from the knowledge base randomly without any criteria.
  • Question Similarity Selection [33] retrieves examples by selecting candidate questions with the highest similarity to the target question.
  • DAIL Selection [10] retrieves examples by jointly considering question- and query-level similarity. Domain-specific tokens in questions are first masked, after which candidates are ranked by question similarity and filtered based on whether their query Jaccard similarity to a pre-generated reference SQL exceeds a predefined threshold.
  • Question-SQL Similarity Selection encodes the target question and each candidate SQL into a shared embedding space with the untrained model, and selects the examples with the highest embedding similarity.
  • PreSQL-SQL Similarity Selection ranks examples by the embedding similarity between each candidate SQL and a pre-generated SQL for the target question, and selects the most similar ones.
  • Question-AST Similarity Selection parses each candidate SQL into an abstract syntax tree (AST) with SQLGlot and converts it into a compact, structure-preserving representation, which keeps all child expressions to retain query structure, while pruning semantically redundant fields (e.g., None arguments and default flags). Candidates are then ranked by the embedding similarity between the target question and the AST representation.
TABLE III: Performance comparison of retrieval strategies. “Upper limit”: DAIL selection retrieval using ground-truth SQL. “API” indicates whether LLM API calls are required during retrieval, with an average token cost of 517 tokens per query.
Retrieval Strategy API BIRD Spider Spider Spider- Spider- Spider- Average
dev dev test DK Realistic Syn
Reference strategies without mask
Zero-shot - 52.5 74.1 75.5 60.2 71.1 62.2 65.9
Random selection 56.6 78.3 79.8 64.5 72.2 66.6 69.7
Question similarity 58.3 79.6 81.9 66.2 74.4 67.7 71.4
DAIL selection 58.8 80.0 82.2 66.4 75.2 68.9 71.9
PreSQL-SQL similarity 59.1 78.9 81.0 63.7 74.2 69.6 71.1
Question-AST similarity 58.2 79.6 80.7 64.5 73.6 65.7 70.4
Question-SQL similarity 58.7 80.9 81.7 65.6 74.2 66.4 71.3
Upper limit 61.8 83.9 85.7 70.8 77.4 72.7 75.4
SQLFlow-part+Spider+BIRD AST 59.1 80.7 81.9 66.4 75.4 67.7 71.9
SQLFlow-part+Spider+BIRD SQL 59.3 81.0 82.1 67.3 75.6 70.1 72.6
Reference strategies with mask
Question similarity 59.0 81.0 83.1 67.5 75.6 69.9 72.7
DAIL selection 59.4 80.8 82.9 67.1 75.2 69.8 72.5
PreSQL-SQL similarity 59.3 77.5 78.5 63.6 71.9 67.5 69.7
Question-AST similarity 56.8 78.2 79.6 65.4 73.0 67.2 70.0
Question-SQL similarity 57.8 80.1 80.3 65.6 73.4 69.0 71.0
Upper limit 61.6 83.8 85.1 69.2 78.7 73.6 75.3
SQLFlow-part+Spider+BIRD AST 59.5 81.3 83.0 68.8 76.4 70.9 73.3
SQLFlow-part+Spider+BIRD SQL 61.0 81.2 83.5 67.7 77.2 71.1 73.6
IV-A5 Implementation Details
@@
To validate the effectiveness of our framework, we conduct full-parameter SFT on two widely adopted open-source LLMs: Meta-Llama-3.1-8B-Instruct [13] and Qwen2.5-Coder-7B-Instruct [20]. We use the entire SQLFlow dataset as training data. The training configuration includes a learning rate of $2\times 10^{-5}$, 2 training epochs, and a warmup ratio of 15%. During LLM inference, we explore two decoding strategies: greedy decoding and majority voting. Greedy decoding (denoted as Gre in tables) uses a temperature of 0 to produce deterministic outputs. Majority voting (denoted as Maj in tables) samples 8 candidate responses per input at a temperature of 0.8. From these candidates, valid SQL queries are extracted and executed; the final prediction is selected as the query whose execution result receives the most votes.
@@
As for the retrieval module, we fine-tune a Qwen3-Embedding-0.6B model [55] via SFT. To construct the training data, we use a filtered subset of the SQLFlow-Spider and SQLFlow-BIRD datasets (described in Section III-B7), retaining only SQL queries that begin with a SELECT clause. This process yields 67,570 training instances, which we refer to as SQLFlow-part. The embedding model is trained for three epochs with a learning rate of $6\times 10^{-6}$ and a per-device batch size of 2. During training, the NL question is treated as the query, and its corresponding SQL query is used as the positive example. For each query, 30 negative SQL examples are randomly sampled from the remaining dataset. At inference time, we compute question–SQL similarity using cosine similarity between embeddings generated by both the original and the fine-tuned embedding models. We apply the masking strategy proposed in DAIL–SQL [10] during both training and inference. Retrieved examples are incorporated into the prompt using a 5-shot setting for LLM inference. For experiments involving closed-source LLMs, we use GPT-4o with greedy decoding and set the temperature to zero. Prompt construction follows the organization strategy introduced in DAIL. In addition to SQLFlow-part, we construct two baseline datasets for comparison. The first baseline combines the standard Spider-train and BIRD-train splits. The second baseline, denoted as SynSQL-part, consists of 67,570 SQL queries randomly sampled from SynSQL-2.5M [29], matched in size with SQLFlow-part to ensure a fair comparison.
@@ IV-B Performance of LLM Fine-tuning
As shown in Table II, the data generated by our framework leads to consistent performance improvements across multiple mainstream benchmarks, demonstrating the effectiveness of Text2SQL-Flow. We fine-tune two widely used open-source models, Meta-Llama-3.1-8B-Instruct and Qwen2.5-Coder-7B-Instruct, to evaluate the general applicability of our data. In both cases, models trained on our generated data significantly outperformed their respective baselines as well as other models. When fine-tuned with SQLFlow data, Qwen2.5-Coder-7B-Instruct achieved notable gains: execution accuracy (“Gre”) on Spider-dev increased from 80.7% to 83.4% (+2.7%), on BIRD-dev from 53.2% to 56.8% (+3.6%), and on the challenging EHRSQL benchmark from 34.6% to 53.7% (+19.1%). These results confirm two key points: (1) SFT substantially enhances model performance on Text-to-SQL tasks, and (2) the data generated via our Text2SQL-Flow framework exhibits high quality and strong training utility. The consistent improvements across two distinct model architectures demonstrate the broad effectiveness of our dataset in boosting model performance.
@@
In comparison with other training datasets, our data demonstrates clear advantages. Specifically, models trained with SQLFlow (75K) not only outperform the baseline models but also consistently surpass SynSQL (75K), highlighting the robustness of our data generation approach. Remarkably, despite being trained on a substantially smaller dataset, the model fine-tuned with SQLFlow (75K) achieves performance comparable to SynSQL-2.5M on several challenging benchmarks. Furthermore, when real-world seed data, including Spider-train, BIRD-train, and EHRSQL-train, are incorporated into the training set, the resulting model attains state-of-the-art (SOTA) performance across multiple benchmarks. These results indicate that expanding seed data with SQLFlow effectively enhances model performance.
TABLE IV: Comparison of training data constructed under different ablation settings. # Tokens per SQL indicates the average tokens consumed to generate each augmented SQL.
Training Setting Spider Spider BIRD # Token
dev test dev per SQL
Baseline 78.3 79.8 52.5 -
Full Pipeline 81.7 83.2 55.5 4909
Auxiliary Database Selection
w/o Auxiliary Database 81.1 82.3 54.8 0
Random Selection 80.9 81.7 54.0
SQL Augmentation Strategy
w/o DVT 80.9 82.5 54.0 857
w/o QSM 80.1 81.8 53.4
w/o BLC 79.7 82.3 54.0
w/o CE 79.7 81.9 53.8
w/o ASF 81.1 81.5 55.5
w/o PO 79.9 81.7 54.2
w/o Executability Filter 80.1 82.3 54.1 0
Question Style
w/o TF 80.4 83.1 53.5 1742
w/o SSI 80.1 82.0 55.4
w/o IDC 80.9 82.3 54.0
w/o IP 81.5 82.8 54.0
w/o Correspondence Filter 80.5 80.0 53.4 709
w/o Prompt Generation - - - 0
w/o Chain-of-Thought 75.5 78.1 53.3 1601
IV-C Ablation Study of the Text2SQL-Flow Pipeline
We conduct an ablation study to analyze the contribution of each component in Text2SQL-Flow. Results are reported in Table IV. We select 20 databases from Spider as original databases and 20 databases from BIRD as auxiliary databases, from which 200 representative SQLs are sampled per source using SQL-structure-based bucketing, resulting in 400 seed SQLs. Each seed SQL is augmented using all six SQL augmentation strategies, each instantiated twice, yielding $4,800$ augmented SQLs. Each augmented SQL is paired with four question styles, producing 19,20019{,}200 SQL–question pairs. All components of our pipeline are enabled unless explicitly ablated. The Baseline (line 1) is trained solely on the seed SQLs, whereas all other settings are trained on augmented data with a fixed size of 2,400 instances. The Full Pipeline (line 2) samples training data evenly from Spider and BIRD. w/o Auxiliary Database (line 3) draws samples exclusively from the Spider source while keeping the total number of sampled instances fixed at 2,400, whereas Random Database Selection (line 4) replaces the structured auxiliary database selection strategy with random sampling. For SQL augmentation and question-style ablations, one strategy or style is removed at a time, with uniform sampling over the remaining ones; the corresponding category names follow the abbreviations defined in Section III-B1 and Section III-B3. In w/o Executability Filter (line 11) and w/o Correspondence Filter (line 16), invalid or misaligned samples are retained during training. w/o Chain-of-Thought (line 18) removes the CoT rationales while keeping the same training samples.
Full Pipeline consistently outperforms the baseline, demonstrating the effectiveness of Text2SQL-Flow. Removing auxiliary databases degrades performance, while random database selection further harms results, validating the necessity of auxiliary databases and our structured coverage strategy. Ablating any SQL augmentation strategy or question style leads to performance drops, indicating that both SQL-level diversity and linguistic variation are important. Disabling executability or alignment filters consistently hurts performance, highlighting the importance of data quality control. Removing CoT supervision causes the most severe degradation, underscoring the critical role of explicit reasoning signals.
TABLE V: Performance of models trained on the same amount of synthetic data generated by different LLMs. “API cost” is reported in USD per million tokens and is set to zero for open-source models, excluding local compute overhead.
LLM Backbone API Cost Spider dev Spider test BIRD dev
Baseline - 78.7 78.0 50.3
Closed-source LLMs
Claude-3.7-Sonnet 12 80.8 83.3 50.9
Gemini-2.5-Flash 0.6 82.8 83.0 52.3
GPT-3.5-Turbo 0.5 80.9 81.7 51.6
GPT-4o 2.5 81.2 81.7 54.1
Open-source LLMs
DeepSeek-V3 0 79.8 81.0 50.7
GLM-4.6 0 80.9 81.0 51.4
Qwen-3-235B 0 80.5 80.6 50.8
IV-D Robustness to the Choice of Data Generation Models
Many components of Text2SQL-Flow rely on LLMs, making the choice of the data generation backbone a potentially influential factor for the quality of the augmented data. To examine this sensitivity, we instantiate the same SQL data augmentation pipeline with a diverse set of both closed-source and open-source LLMs. Throughout this study, we keep the overall pipeline design, augmentation procedures, and the scale of synthetic data fixed, and vary only the LLM used for data generation. Specifically, each model is prompted to generate 1,000 synthetic training examples based on 500 samples randomly selected from the Spider training set. The resulting augmented data are then used to train the downstream Text-to-SQL model. Baseline (line 1) refers to the model trained using data constructed solely from these 500 samples.
As shown in Table V, the downstream performance remains competitive across a wide range of data generation models. While stronger LLMs tend to produce slightly better synthetic data, even relatively weaker or low-cost models, including GPT-3.5-Turbo and open-source LLMs, still yield solid performance on benchmarks. This suggests that Text2SQL-Flow does not critically depend on highly sophisticated LLM reasoning during data generation. Instead, its robustness primarily stems from the structured augmentation process, executability filtering, and semantic alignment, which together mitigate noise introduced by imperfect generation models. Across all LLM backbones, the performance variation remains within a moderate range, with all models outperforming the baseline, without altering the overall effectiveness of the pipeline.
IV-E Comparison of Retrieval Strategies